A plate previously cold worked at 45% is 50mm thick.Its is then cold worked further by 27mm. What is the percentage cold work?
1
Expert's answer
2020-12-07T03:48:23-0500

Calculate thickness before 45% cold is as follows

tinitial "= \frac{50}{0.55}"


= 90.90 mm

Calculate the cold worked percent as follows

% cold-worked = "\frac{tinital - tfinal}{tinitial}"

= "\frac{90.90-(50-27)}{90.90}\times 100"


= "\frac{67.9}{90.90}\times 100"


= 74.6%